OpenStudy (anonymous):

A circle has center (0,0) and radius 6. a)write the equation of the circle. b) the point b (4,k) is on the circle. Determine the value of k.

OpenStudy (anonymous):

The general equation of a circle is \[(x-h)^2+(y-k)^2=r^2 \] So a is \[x^2+y^2=r^2 \]

OpenStudy (anonymous):

@eaiufhea Can you follow purplec16's guide?